OpenStudy (anonymous):

MATH HELP?! GRADUATION OF 10TH GRADE ON THE LINE! #1 A car’s headlight contains a parabolic reflector. A special bulb with two filaments is used to produce the high and low beams. The filament placed at the focus produces the high beam and the filament placed off-focus produces the low beam. The equation of the reflector is y=1/12 x^2. How far from the vertex should the filament for the high beam be placed? The x and y values are in centimeters. #2 Concentric circles have the same center but different radii. Write equations for two concentric circles centered at ( -3, 5) with radii of 7 and 9.

OpenStudy (mathmale):

First: I believe you can completely ignore all the info about the low beam filament: A car’s headlight contains a parabolic reflector. A special bulb with two filaments is used to produce the high and low beams. The filament placed at the focus produces the high beam. The equation of the reflector is y=1/12 x^2. How far from the vertex should the filament for the high beam be placed? The x and y values are in centimeters. Hint: The appropriate equation of a parabola that relates to the distance between the focus and the vertex is \[4py=x^2\] Compare this equation to your \[y=\frac{ 1 }{ 12 }x^2\] and determine the value of p. this value of p is the answer to, "How far from the vertex should the filament for the high beam be placed?"

OpenStudy (anonymous):

No it isnt an exam sir, it is homework help.

OpenStudy (mathmale):

@kitties33: In no way did I mean to imply that you were asking for help with a test question. don't worry about that!

OpenStudy (anonymous):

No there was another person who had posted :b Okay I still dont get it though :/

OpenStudy (mathmale):

Rewrite y = (1/12)x^2 as 12y=x^2.

OpenStudy (mathmale):

Now please compare this 12y=x^2 to 4py=x^2 and solve for p.

OpenStudy (anonymous):

I am feeling extremely dumb because how do you compare it? is p=12 or does it = 3

OpenStudy (anonymous):

I think it is 3 because you multiply it to the 4 for the 12 correct?

OpenStudy (mathmale):

12y=x^2=4py, so, 12y =4py (divide both sides by y) 4p = 12 p = 12/4 = 3. YES!

OpenStudy (mathmale):

So: how far should we place the high beam filament from the vertex of this parabolic reflector?

OpenStudy (anonymous):

3 centimeters? o.O

OpenStudy (mathmale):

Perfect. Nice work!!

OpenStudy (anonymous):

Thank you !!!

OpenStudy (mathmale):

Hope to work with you again. For now: I have to get off the Internet!

OpenStudy (anonymous):

Alright thank you for the help :)